OpenStudy (anonymous):

CHECK MY ANSWER! What is the mean absolute deviation (MAD) of the data set 2,2,5? A. 0 B. 2/3 C. 4/3****** D. 1/3

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

im pretty sure its C cuz i found out the mean and its 3; 2+2+5/3=3 and then I2-3I+I2+3I+I5-3I=4 and then divide by the mean which is 3 so its 4/3... am i correct?

OpenStudy (anonymous):

@iambatman @EclipsedStar @e.mccormick

OpenStudy (anonymous):

@pinklion23

OpenStudy (pinklion23):

im here sorry i was eating :)

OpenStudy (anonymous):

its ok :) is it right?

OpenStudy (pinklion23):

no D is corecct

OpenStudy (anonymous):

how?

OpenStudy (pinklion23):

okay so 2+2+5=?

OpenStudy (anonymous):

9

OpenStudy (pinklion23):

yep now 9/how many numbers are in your data set?

OpenStudy (anonymous):

3 9/3=3=mean

OpenStudy (pinklion23):

yep

OpenStudy (anonymous):

I2-3I+I2+3I+I5-3I=4 and then divide by the mean which is 3 so its 4/3.

OpenStudy (pinklion23):

so your numbers in your data set are how many units away from your mean? for instance if your mean is 5 and one of your numbers is 7 then it is 2 units away from 5

OpenStudy (anonymous):

i already took the quiz and C was correct...

OpenStudy (pinklion23):

oh :/

OpenStudy (pinklion23):

well then how did you get 3/4?

OpenStudy (anonymous):

i explained it i just wanted to see if i was correct and i was :D

OpenStudy (anonymous):

its 4/3

OpenStudy (pinklion23):

ho i see you did a different thing sorry

OpenStudy (pinklion23):

i did the next step which is 4/3 which is 1/3

OpenStudy (pinklion23):

thx for the medal anyway :)

OpenStudy (anonymous):

its ok :) lol 1/3 = 0.33

OpenStudy (pinklion23):

well i did not go that far i only did 4/3=1/3 :)

OpenStudy (anonymous):

your welcome :D

OpenStudy (anonymous):

4/3 = 1.33 not 0.33 which = 1/3

OpenStudy (anonymous):

its fine i already submitted it so who cares xD lol

OpenStudy (pinklion23):

okay have any other questions?